# Latent Adaptive Planner for Dynamic Manipulation

### v1

@@ Abstract
This paper presents Latent Adaptive Planner (LAP), a novel approach for dynamic nonprehensile manipulation tasks that formulates planning as latent space inference, effectively learned from human demonstration videos. Our method addresses key challenges in visuomotor policy learning through a principled variational replanning framework that maintains temporal consistency while efficiently adapting to environmental changes. LAP employs Bayesian updating in latent space to incrementally refine plans as new observations become available, striking an optimal balance between computational efficiency and real-time adaptability. We bridge the embodiment gap between humans and robots through model-based proportional mapping that regenerates accurate kinematic-dynamic joint states and object positions from human demonstrations. Experimental evaluations across multiple complex manipulation benchmarks demonstrate that LAP achieves state-of-the-art performance, outperforming existing approaches in success rate, trajectory smoothness, and energy efficiency, particularly in dynamic adaptation scenarios. Our approach enables robots to perform complex interactions with human-like adaptability while providing an expandable framework applicable to diverse robotic platforms using the same human demonstration videos.
Keywords: Imitation Learning, Robotics, Dynamic Nonprehensile Manipulation, Latent Space Planning, Classical Variational Bayes, Video-based Skill Acquisition
@@ 1 Introduction
Dynamic manipulation represents a sophisticated robotic capability that involves controlling objects through rapid changes in contact states, inertial forces, impacts, and complex physical interactions [1, 2]. While everyday human actions such as throwing, catching, and rapid object transfers all rely on dynamic physical interactions, current robotics research has predominantly focused on static or quasi-static manipulation [3, 4]. This static manipulation paradigm inherently limits robots to slow speeds, conservative grasping forces with high safety factors, and restricted motion ranges. Achieving true dynamic manipulation requires real-time consideration of diverse physical properties of objects (mass, friction, elasticity) [5, 6], changes in contact states, and joint torques [7, 8, 9], yet these dynamic characteristics are difficult to model analytically and vary significantly across different objects [10, 11].
The robotics community has attempted to address these complex dynamic manipulation challenges through two primary approaches. First, physics-based modeling approaches have sought to mathematically represent dynamics precisely, but perfectly modeling real-world uncertainties and diverse object properties remains practically impossible [12]. Second, machine learning-based approaches have attempted to learn system dynamics from data, but most consider only end-effector positions while overlooking critical dynamic elements such as contact information, joint torques, and grasping forces [3, 4, 13, 14]. While learning-based approaches often utilize data collected through teleoperation, this method faces significant practical limitations: high implementation costs, logistical challenges for achieving data scale and diversity, and technical difficulties in effectively capturing complete physical interaction data [15, 16]. Although advanced teleoperation systems capable of measuring and transmitting grasping forces, contact states, and joint torques exist in specialized research settings [17], they remain insufficiently deployed for large-scale data collection. These limitations present substantial barriers to implementing human-level adaptive and efficient dynamic manipulation capabilities in robots [18].
Traditional approaches often struggle with the temporal consistency and trajectory smoothness needed for natural robotic movements, particularly when transferring skills from human demonstrations to robotic systems [19]. While recent advances in offline reinforcement learning and imitation learning have shown promise, they often fail to generalize to scenarios requiring real-time adaptation. The emergence of generative modeling techniques for sequential decision-making has opened new avenues for addressing these challenges such as sequence modeling [20] and diffusion models [3]. However, these methods exhibit limitations in handling the long-term dependencies and environmental contingencies inherent in dynamic manipulation tasks.
@@
In this work, we introduce Latent Adaptive Planner (LAP), which formulates planning as latent space inference effectively learned from human demonstration videos. Our approach addresses key challenges in visuomotor policy learning through a principled variational replanning framework that maintains temporal consistency while efficiently adapting to environmental changes. By leveraging a latent variable that functions as an abstract plan, LAP enables more coherent long-horizon planning while maintaining the flexibility needed for real-time adaptation. The system employs Bayesian updating in latent space to incrementally refine plans as new observations become available, striking an optimal balance between computational efficiency and real-time adaptability. Furthermore, by utilizing model-based proportional mapping to regenerate accurate kinematic-dynamic joint states and object positions from human demonstration videos, our approach bridges the gap between human motion and robotic execution, capturing the adaptive and energy-efficient qualities inherent in human manipulation skills. Importantly, this regeneration methodology allows us to create versatile datasets suitable for training diverse robotic platforms on the same manipulation tasks without requiring direct robot data collection.
@@ 2 Related Works
Imitation learning enables robots to acquire skills by mimicking demonstrations, transforming complex control into supervised learning problems that map observations to actions [35], though traditional approaches like Behavior Cloning struggle with distribution shift during deployment [36]. Recent approaches have reframed decision-making as sequence modeling [20, 37] to capture temporal dependencies in state-action sequences. Diffusion models offer an alternative paradigm for imitation learning, with Chi et al. [3] generating multimodal action distributions through iterative denoising processes and Janner et al. [38] leveraging classifier-guided diffusion models.
@@ 3 Method
3.1 Robot Model Based Data Regeneration from Human Demonstration Videos
@@ 3.2 Latent Adaptive Planner (LAP)
Given a trajectory ${\mathbf{x}}$ consisting of observations and actions from demonstrations, our LAP defines a joint probability distribution:
@@
where ${\mathbf{z}}\in\mathbb{R}^{d}$ is the latent vector representing the plan, $p({\mathbf{z}})$ is the prior model, and $p_{\theta}({\mathbf{x}}|{\mathbf{z}})$ is the trajectory generator. The prior model $p({\mathbf{z}})$ is defined as an isotropic Gaussian ${\mathbf{z}}\sim\mathcal{N}(0,I_{d})$.
The trajectory generator $p_{\theta}({\mathbf{x}}|{\mathbf{z}})$ is a conditional autoregressive model that produces actions based on the current state, historical context, and the latent plan:
@@
where ${\mathbf{x}}_{(t)}=(o_{t},a_{t})$, and $K$ is the context window length. The observation $o_{t}$ includes box position, contact state, and previous timestep joint positions, velocities, and torques ($\mathbf{q}_{t-1}$, $\dot{\mathbf{q}}_{t-1}$, $\bm{\tau}_{t-1}$). Action $a_{t}$ consists of current timestep joint positions, velocities, and torques ($\mathbf{q}_{t}$, $\dot{\mathbf{q}}_{t}$, $\bm{\tau}_{t}$). ${\mathbf{x}}_{(0)}$ is a special learnable token. $\theta$ is implemented using a causal Transformer model [47]. The latent plan vector ${\mathbf{z}}$ controls each step of action generation through cross-attention mechanisms as $p_{\theta}(a_{t}|o_{t-K:t-1},a_{t-K:t-1},{\mathbf{z}})$. The action is assumed to follow a single-mode Gaussian distribution with fixed variance, i.e. $a_{t}\sim\mathcal{N}(g_{\theta}(o_{t-K:t},a_{t-K:t-1},{\mathbf{z}}),I_{|a|})$.
@@
Variational Replanning
We propose a novel approach termed variational replanning that implements a principled Bayesian update procedure in the latent space, realizing the concept of planning as latent space inference [41], which significantly improves the model efficiency.
@@
The essence of our approach is treating the previously inferred distribution $q({\mathbf{z}}|{\mathbf{x}}_{0:t})$ as the prior belief and updating it with a small number of gradient steps to obtain $q({\mathbf{z}}|{\mathbf{x}}_{0:t+\Delta})$. As $t+\Delta\rightarrow T$ approaches the full trajectory length, the variational posterior $\mathcal{N}(\bm{\mu}_{t+\Delta},\bm{\sigma}^{2}_{t+\Delta})$ converges to the global posterior $\mathcal{N}(\bm{\mu}_{T},\bm{\sigma}^{2}_{T})$ that would be inferred given the complete trajectory. This establishes theoretical consistency between our incremental replanning approach and the optimal plan that would be determined with complete information. Detailed algorithms can be found in Appendix E. The integrated pipeline can be found in Figure 2.
@@ 4 Experimental Results
We evaluate our LAP on a nonprehensile dynamic manipulation task of box receiving, comparing its performance against several baseline methods including rule-based planning, behavior cloning, and Diffusion Policy approaches. This task represents a challenging example of nonprehensile dynamic manipulation, where the robot must initially interact with and control objects without stable grasping, using the dynamics of the arm movement to stabilize and then secure the incoming object.
@@ 4.1 Experimental Setup
Box receiving constitutes the primary focus of this study, a challenging nonprehensile dynamic manipulation task where the robotic system must intercept, temporarily control, and securely grasp boxes of varying sizes, weights, and shapes handed or tossed by human operators. This task requires real-time trajectory adaptation, impact anticipation, and force modulation as the robot first manages dynamics using arm surfaces before transitioning to a stable grasp.
@@
Figure 3: Impact-aware retreat trajectory for box receiving motion learned with the Latent Adaptive Planner (LAP). The left panel shows a human demonstration where the subject absorbs impact by yielding their arm along the trajectory of the incoming box before returning to the nominal pose. The right panel shows the robot reproducing this compliant motion using LAP. The red curves indicate the impact-aware retreat trajectory that minimizes energy consumption during dynamic interaction.
@@
Robot Method Success (%) Energy (J) Success (%) Energy (J) Success (%) Energy (J)
@@
[…]
@@
[…]
@@
[…]
@@ 4.2 Results and Analysis
Our experiments demonstrate that LAP consistently outperforms the baselines in all metrics and box types on both robot platforms. Table 1 summarizes the quantitative results for Robot A and Robot B, respectively. Figure 3 visualizes the impact-aware retreat trajectory for the the box receiving motion learned by LAP using Robot B.
@@ 5 Conclusion
Our Latent Adaptive Planner (LAP) demonstrates significant advantages in nonprehensile dynamic manipulation tasks, achieving higher success rates while consuming less energy and maintaining lower peak torques compared to existing approaches. By introducing a latent plan as a temporal abstraction of the trajectory, our method achieves consistent long-horizon behavior while maintaining the flexibility required for dynamic environments.
The performance improvements observed across different box types highlight the effectiveness of our adaptive planning approach. Furthermore, our experiments with both Robot A and Robot B confirm that LAP successfully transfers to different robot configurations. By regenerating appropriate training data from the same human demonstration videos for each platform, our method effectively adapts to the distinct kinematic and dynamic properties of different robots. This cross-platform success validates the generalizability of our approach and demonstrates that the combination of data regeneration and latent planning provides a robust framework for dynamic manipulation tasks regardless of the specific robot embodiment.
@@
Despite the promising results demonstrated by our Latent Adaptive Planner, several limitations warrant acknowledgment and present opportunities for future research:
First, our approach employs a simplified binary contact representation that does not capture the continuous nature of physical interactions. More critically, we approximate contact locations by assuming the most significant impact occurs at the wrist position, rather than precisely estimating the actual contact points and force distributions. This simplification fundamentally limits our ability to model the complex dynamics of manipulation tasks where contact may occur across multiple points with varying intensities and directions.
Second, to reduce kinematic complexity, we restricted our human demonstration mapping to three primary degrees of freedom (shoulder, elbow, and wrist pitch), rather than utilizing the full seven degrees of freedom available in human arm movements. This simplification excludes potentially more efficient motion patterns that would be possible with base mobility and waist joint articulation in a full humanoid configuration. By constraining our system to a fixed-base manipulator paradigm, we limit the natural whole-body coordination that humans employ during dynamic manipulation tasks, where weight shifting and torso movements significantly contribute to task performance, energy efficiency, and stability during dynamic interactions.
Third, our reliance on 2D object and human pose estimation introduces inherent limitations in spatial reasoning, as it cannot fully capture depth-dependent interactions or account for occlusions that occur during dynamic movements. This dimensional reduction potentially impacts the fidelity of learned manipulation strategies, especially in tasks requiring precise depth perception.
Addressing these limitations will require advances in both modeling approaches and sensing technologies to capture more detailed contact dynamics, improved dimensional representation, and expanded degrees of freedom in robotic execution systems.
From a generative modeling perspective, the Latent Adaptive Planner relies on a relatively simple prior model (isotropic Gaussian) that may not adequately capture the complex, multi-modal nature of expert demonstration trajectories. More structured prior models [50, 51, 52, 53] could potentially provide richer representations of the latent space, enabling more nuanced planning capabilities and improved generalization to novel scenarios. Future work will explore these advanced modeling approaches to enhance the representational capacity and performance of our framework.

### v2

@@ Abstract
We present the Latent Adaptive Planner (LAP), a trajectory‑level latent‑variable policy for dynamic nonprehensile manipulation (e.g., box catching) that formulates planning as inference in a low‑dimensional latent space and is learned effectively from human demonstration videos. During execution, LAP achieves real-time adaptation by maintaining a posterior over the latent plan and performing variational replanning as new observations arrive. To bridge the embodiment gap between humans and robots, we introduce a model-based proportional mapping that regenerates accurate kinematic-dynamic joint states and object positions from human demonstrations. Through challenging box catching experiments with varying object properties, LAP demonstrates superior success rates, trajectory smoothness, and energy efficiency by learning human-like compliant motions and adaptive behaviors. Overall, LAP enables dynamic manipulation with real-time adaptation and successfully transfer accross heterogeneous robot platforms using the same human demonstration videos.
Keywords: Imitation Learning, Dynamic Nonprehensile Manipulation, Latent Space Planning, Classical Variational Bayes, Test-time Adaptation
@@ 1 Introduction
Dynamic manipulation, which involves controlling objects through rapid contact changes and complex physical interactions [1, 2], remains a fundamental challenge in robotics. While humans naturally perform throwing, catching, and rapid transfers, robots remain confined to slow, conservative movements. Achieving true dynamic manipulation requires real-time consideration of diverse physical properties of objects (mass, friction, elasticity) [3, 4], changes in contact states, and joint torques [5, 6, 7], yet these dynamic characteristics are difficult to model analytically and vary significantly across different objects [8, 9]. Recent advances like Tossingbot [10], FlingBot [11], and Chi et al. [12] have shown progress in specific scenarios, yet predominantly rely on self-supervised learning requiring thousands of robot trials. For this reason, this approach lacks scalability since exploration with heavier objects or mobile-based platforms like humanoids become unsafe and leads to significant hardware damage upon failure. While simulation offers safer training, accurately modeling dynamic manipulation physics remains challenging, limiting sim-to-real transfer.
Recently, imitation learning has shown promising results on static or quasi-static manipulation tasks. However, recent approaches such as diffusion policy are limited by slow inference speeds that cannot meet the real-time requirements of dynamic manipulation. Moreover, most datasets for policy training consider only end-effector positions while overlooking critical dynamic elements such as contact information, joint torques, and grasping forces [13, 14, 15]. While more accurate data can be collected through teleoperation, this approach presents practical barriers including high implementation costs, logistical challenges for achieving scale and diversity, and technical difficulties in capturing complete physical interaction data [16, 17, 18]. Additionally, demonstration-based learning approaches struggle with temporal consistency and trajectory smoothness when transferring skills from human demonstrations to robotic systems [19]. Although recent generative modeling approaches such as sequence modeling [20] and diffusion models [13] have addressed some of these challenges, they still exhibit limitations in handling the long-term dependencies and environmental contingencies inherent in dynamic manipulation tasks [18].
@@
In this work, we introduce Latent Adaptive Planner (LAP), which formulates planning as latent space inference. Our approach addresses key challenges in visuomotor policy learning through a principled variational replanning framework that maintains temporal consistency while efficiently adapting to environmental changes. By leveraging a latent variable that functions as an abstract plan, LAP enables more coherent long-horizon planning while maintaining the flexibility needed for real-time adaptation. The system employs Bayesian updating in latent space to incrementally refine plans as new observations become available, balancing computational efficiency with real-time adaptability. Furthermore, by model-based proportional mapping to regenerate accurate kinematic-dynamic joint states and object positions from human demonstration videos, our approach bridges the gap between human motion and robotic execution, capturing the adaptive and energy-efficient qualities inherent in human manipulation skills. Importantly, this regeneration methodology allows us to create versatile datasets suitable for training diverse robotic platforms on the same manipulation tasks without requiring direct robot data collection.
We demonstrate our approach through box catching, one of the most challenging reactive manipulation tasks. Unlike controlled release scenarios such as throwing and fling, catching requires real-time adaptation to unpredictable trajectories while managing objects without secure grasps [21, 22]. Boxes which have non-trivial weight and volumes tumble chaotically with asymmetric drag, forcing robots to predict contact locations and orientations while coordinating dual-arm movements with millisecond-scale precision and determining appropriate contact forces to absorb impact without dropping the object [21, 23]. Through this task, we show that imitation learning can achieve energy-efficient dynamic manipulation while bypassing the intractable formulations required by optimization methods, demonstrating the potential for broader dynamic manipulation challenges.
@@ 2 Related Works
Imitation learning enables robots to acquire skills by mimicking demonstrations, transforming complex control into supervised learning problems that map observations to actions [38], though traditional approaches like Behavior Cloning struggle with distribution shift during deployment [39]. Recent approaches have reframed decision-making as sequence modeling [20, 40] to capture temporal dependencies in state-action sequences. Diffusion models offer an alternative paradigm for imitation learning, with Chi et al. [13] generating multi-modal action distributions through iterative denoising processes and Janner et al. [41] leveraging classifier-guided diffusion models.
@@ 3 Method
3.1 Robot Model-Based Data Regeneration from Human Demonstration Videos
@@ 3.2 Latent Adaptive Planner (LAP)
Given a trajectory ${\mathbf{x}}=\{(o_{t},a_{t})\}_{t=1}^{T}$ consisting of observations and actions from demonstrations, LAP defines a joint probability distribution:
@@
where ${\mathbf{z}}\in\mathbb{R}^{d}$ is the latent plan vector, $p({\mathbf{z}})$ is the prior model, and $p_{\theta}({\mathbf{x}}|{\mathbf{z}})$ is the trajectory generator. The prior model $p({\mathbf{z}})$ is an isotropic Gaussian ${\mathbf{z}}\sim\mathcal{N}(0,I_{d})$.
The trajectory generator $p_{\theta}({\mathbf{x}}|{\mathbf{z}})$ is a conditional autoregressive model that produces actions based on the current state, historical context, and the latent plan vector:
@@
where ${\mathbf{x}}_{(t)}=(o_{t},a_{t})$, and $K$ is the context length. The observation $o_{t}$ includes box position, contact state, and previous timestep joint positions, velocities, and torques ($\mathbf{q}_{t-1}$, $\dot{\mathbf{q}}_{t-1}$, $\bm{\tau}_{t-1}$). Action $a_{t}$ consists of current timestep joint positions, velocities, and torques ($\mathbf{q}_{t}$, $\dot{\mathbf{q}}_{t}$, $\bm{\tau}_{t}$). ${\mathbf{x}}_{(0)}$ is a special learnable token. $\theta$ is implemented using a causal Transformer model [50]. The latent plan vector $\mathbf{z}$} controls each step of action generation through cross-attention mechanisms as $p_{\theta}(a_{t}|o_{t-K:t-1},a_{t-K:t-1},{\mathbf{z}})$. The action is assumed to follow a unimodal Gaussian distribution with fixed variance, i.e. $a_{t}\sim\mathcal{N}(g_{\theta}(o_{t-K:t},a_{t-K:t-1},{\mathbf{z}}),I_{|a|})$.
@@
Variational Replanning for Test-time Adaptation
We propose the variational replanning method that implements a principled Bayesian updating in the latent space, realizing the concept of planning as latent space inference [44], which significantly improves the model efficiency and stability.
@@
The essence of our approach is treating the previously inferred distribution $q({\mathbf{z}}|{\mathbf{x}}_{0:t})$ as the prior belief and updating it with a small number of gradient steps to obtain $q({\mathbf{z}}|{\mathbf{x}}_{0:t+\Delta})$. As $t+\Delta\rightarrow T$ approaches the full trajectory length, the variational posterior $\mathcal{N}(\bm{\mu}_{t+\Delta},\bm{\sigma}^{2}_{t+\Delta})$ converges to the posterior distribution of $\mathcal{N}(\bm{\mu}_{T},\bm{\sigma}^{2}_{T})$ that would be inferred given the complete trajectory. This establishes theoretical consistency between our incremental replanning approach and the optimal plan that would be determined with complete information. Detailed algorithms can be found in Appendix E. The integrated pipeline can be found in Figure 2.
@@ Discussion
Meanwhile, the objective in Equation 11 performs a KL-regularized proximal step around $q_{t}({\mathbf{z}})$: it increases the likelihood of the newly observed segment while staying close to $q_{t}$. Its optimizer is the exponential-tilting update in Equation 10, i.e., $q_{t}$ reweighted by the new-segment likelihood. With $q$ constrained to Gaussians, our single VB step implements a projected version of this tilt, acting as a small “trust-region” move in latent space that yields stable, real-time adaptation.
@@ 4 Experimental Results
We evaluate our LAP on a nonprehensile dynamic manipulation task of box catching, comparing its performance against several baseline methods including rule-based planning, behavior cloning, and Diffusion Policy approaches. This task represents a challenging example of nonprehensile dynamic manipulation, where the robot must initially interact with and control objects without stable grasping, using the dynamics of the arm movement to stabilize and then secure the incoming object.
@@ 4.1 Experimental Setup
Box catching constitutes the primary focus of this study, a challenging nonprehensile dynamic manipulation task where the robotic system must intercept, temporarily control, and securely grasp boxes of varying sizes, weights, and shapes handed or tossed by human operators. This task requires real-time trajectory adaptation, impact anticipation, and force modulation as the robot first manages dynamics using arm surfaces before transitioning to a stable grasp.
@@
Figure 3: Impact-aware retreat trajectory for box catching motion learned with the Latent Adaptive Planner (LAP). The left panel shows a human demonstration where the subject absorbs impact by yielding their arm along the trajectory of the incoming box before returning to the nominal pose. The right panel shows the robot reproducing this compliant motion using LAP. The red curves indicate the impact-aware retreat trajectory that minimizes energy consumption during dynamic interaction.
@@
Robot Method Success (count) Energy (J) Success (count) Energy (J) Success (count) Energy (J)
@@
Robot A Model-based $\mathbf{30/30}$ $74.99$ $\mathbf{30/30}$ $57.31$ $\mathbf{30/30}$ $70.46$
@@
Robot B Model-based $\mathbf{30/30}$ $33.12$ $\mathbf{30/30}$ $33.17$ $\mathbf{30/30}$ $26.09$
BC $27/30$ $15.82$ $28/30$ $15.40$ $\mathbf{30/30}$ $13.78$
@@
LAP (Ours) $\mathbf{30/30}$ $\mathbf{7.14}$ $\mathbf{30/30}$ $\mathbf{6.64}$ $\mathbf{30/30}$ $\mathbf{6.86}$
@@ 4.2 Results and Analysis
Our experiments demonstrate that LAP consistently outperforms the baselines in all metrics and box types on both robot platforms. Table 1 summarizes the quantitative results for Robot A and Robot B, respectively. Figure 3 visualizes the impact-aware retreat trajectory for the the box catching motion learned by LAP using Robot B.
@@ 5 Conclusion
We introduced the Latent Adaptive Planner (LAP), a trajectory-level latent-variable policy that treats planning as inference and is learned from human demonstration videos via a robot-model-based data regeneration pipeline. During execution , LAP performs variational replanning by updating a posterior over the latent plan as new observations arrive, yielding real-time adaptation in dynamic scenes. This test-time adaptation behaves like a small trust-region move in latent space, providing stability without solving a full optimization at every step.
Our experiments demonstrate LAP’s effectiveness across different manipulation scenarios and validate its transferability to diverse robot configurations. By regenerating appropriate training data from the same human demonstration videos for each platform, LAP successfully adapts to distinct kinematic and dynamic properties of different robots. This cross-platform success confirms that the combination of data regeneration and latent planning provides a robust framework for dynamic manipulation tasks regardless of robot embodiment, opening promising directions for scalable visuomotor learning from human demonstrations.
@@
While our Latent Adaptive Planner demonstrates strong performance on dynamic manipulation tasks, several limitations remain and point to clear directions for future work.
Contact Modeling
Our current approach employs a simplified binary contact representation and approximates contact locations at the wrist position. This simplification limits our ability to model the rich dynamics of multi-point contact interactions with varying force distributions, a critical aspect of dexterous manipulation where contact can occur across the entire arm surface with varying force distributions.
Kinematic Constraints
To facilitate human-to-robot mapping, we restricted our demonstration analysis to three primary degrees of freedom (shoulder, elbow, and wrist pitch), excluding the full seven-DOF capability of human arms. This constraint, combined with our fixed-base manipulator assumption, prevents the system from leveraging whole-body coordination strategies that humans naturally employ during dynamic tasks, such as weight shifting and torso movements that contribute significantly to task performance and energy efficiency.
Perception Limitations
Our reliance on 2D pose estimation from monocular video introduces fundamental constraints in spatial reasoning. The lack of depth information limits our ability to accurately reconstruct 3D trajectories and handle occlusions during dynamic movements, potentially affecting the fidelity of learned manipulation strategies in tasks requiring precise spatial awareness.
Prior Model Limitations
From a generative modeling perspective, LAP employs a simple latent prior that may under-represent complex, multi-modal expert behaviors. More structured prior models, such as energy-based models [53, 54, 55, 56], diffusion models [57], and other expressive families [58, 59], could provide richer representations of the latent space, enabling more nuanced planning capabilities and better generalization to novel scenarios.
Future Work
We aim to incorporate continuous contact models with multi-point force estimation, strengthen 3D perception for accurate spatial reasoning, extend to mobile platforms with whole-body coordination, and explore more expressive latent priors. These advances will be important for approaching human-level dexterity while maintaining the real-time performance required for dynamic manipulation.